# A causal random synthetic turbulent velocity field: I. The underlying Gaussian structure

### v1

@@ Abstract
We develop, simulate and extend an initial proposition by Chaves et al. (2003) concerning a random incompressible vector field able to reproduce key ingredients of three-dimensional turbulence in both space and time. In this first article, we focus on the important underlying Gaussian structure that will be generalized in a second article to account for higher-order statistics. Presently, the statistical spatial structure of this velocity field is consistent with a divergence-free fractional Gaussian vector field that encodes all known properties of homogeneous and isotropic fluid turbulence at a given finite Reynolds number, up to second-order statistics. This includes the velocity power spectral density (PSD), with a given regularization at low wave-vectors allowing to introduce the integral length scale in the picture, and an exponential-like dissipative behavior taking place around a viscous-dependent appropriate length scale. The temporal structure of the velocity field is introduced through a stochastic evolution of the respective Fourier modes. In the simplest picture, Fourier modes evolve according to an Ornstein-Uhlenbeck process, where the characteristic time scale depends on the wave-vector amplitude. For consistency with direct numerical simulations (DNSs) of the Navier-Stokes equations, this time scale is inversely proportional to the wave vector amplitude. As a consequence, the characteristic velocity that governs the eddies is independent of their size and is related to the velocity standard deviation, which is consistent with some features of the so-called sweeping effect. To ensure differentiability in time while respecting the causal nature of the evolution, we use the methodology developed by Viggiano et al. (2020) to propose a fully consistent stochastic picture, predicting in particular proper temporal covariance of the Fourier modes. We finally derive analytically all statistical quantities in a continuous setup and develop precise and efficient numerical schemes of the corresponding periodic framework. Both exact predictions and numerical estimations of the model are compared to DNSs provided by the Johns Hopkins database.
@@ 1 Introduction
Fully developed fluid turbulence is emblematic of fluid mechanics and out-of-equilibrium statistical physics. It is usually observed in various natural flows and studied, in simpler situations than those of geophysical flows, in laboratory experiments and numerical simulations of the forced Navier-Stokes equations. An ancient tradition of measurements and observations has allowed the development of a very precise and self-consistent phenomenology capable of describing its statistical structure with few free parameters (Monin & Yaglom, 1971; Tennekes & Lumley, 1972; Frisch, 1995; Pope, 2000). From the theoretical side, the velocity of a turbulent fluid can be seen as the vanishing viscosity limit of the statistically stationary, isotropic, and homogeneous solution of the incompressible Navier-Stokes equations sustained by a random (Eswaran & Pope, 1988), possibly deterministic (Vincent & Meneguzzi, 1991), force. Because of the interplay between non-linear and non-local aspects of this dynamics, with non-trivial influence of viscosity, the rigorous study of these solutions remains a fantastic mathematical challenge. The main purpose of this article is the presentation, extension, and simulation of a spatio-temporal turbulent velocity field that evolves according to a Markovian dynamics in the simplest situation, to a causal one in a more sophisticated version, able to reproduce several key statistical signatures. This stochastic model encompasses the spatial multi-scale nature of the observed turbulent velocity field, and a random evolution in time, which will be shown to obey a causal dynamics. In the infinite Reynolds number limit and when the analysis is restricted to second-order statistical laws, as captured in particular by a Gaussian framework, and assuming the model to evolve according to a Markovian dynamics, the present approach becomes equivalent to that proposed by Chaves et al. (2003). As far as we know, such models are unrelated to Navier-Stokes solutions, and we will therefore consider them as being synthetic, although they reflect important aspects of fluid turbulence.
@@
The purpose of the present article is to propose a random incompressible field that evolves according to a causal dynamics consistent with the temporal structure of turbulence, which has been repeatedly observed, mostly in DNSs (Kaneda et al., 1999; Chevillard et al., 2005; Favier et al., 2010; Canet et al., 2017; Gorbunova et al., 2021). It is more difficult to observe this phenomenon experimental measurements because it requires the absence of a mean flow, which is present in wind tunnels and jets. However, Gorce & Falcon (2022) have conducted measurements of the temporal evolution of a flow stirred by magnetic particles able to generate turbulence without a mean flow, and temporal statistics were found to be similar to those observed in aforementioned numerical simulations. The temporal spectrum, defined as the variance of the temporal Fourier modes at a given frequency $\omega$, behaves in a universal fashion as a power-law $\omega^{-5/3}$”, with thus a similar exponent as observed on the spatial spectrum. Although understood and predicted by Tennekes (1975), this power-law decay can be seen as counterintuitive and differs drastically from similarity arguments which would predict a much faster decrease, i.e. with a 2-2 power-law exponent, as has been observed in the Lagrangian framework, i.e. following fluid particles along their trajectory (Yeung, 2002; Toschi & Bodenschatz, 2009; Pinton & Sawford, 2012). Temporal regularity of the velocity field is imposed by this power-law, which is thus the same as the spatial one, and turns out to have important consequences. In fact, the characteristic velocity of eddies of a given size is independent of that scale and turns out to be the one at large scale: this is the so-called sweeping effect. From a statistical viewpoint, both this power-law decay of the temporal PSD and the scale-independent characteristic velocity of eddies can be included in the model by setting a decorrelation time scale $T_{\mathbf{k}}$ of Fourier modes proportional to the inverse wave-vector amplitude and whose multiplicative constant is inversely proportional to the velocity standard deviation allowing a consistent physical picture and proper dimension. This time scale dependent on $k$ is at the core of the random picture of Chaves et al. (2003).
The proposed random velocity field of Chaves et al. (2003), see also earlier propositions of Komorowski & Papanicolaou (1997) and Fannjiang & Komorowski (2000), encodes all the aforementioned spatial and temporal statistical structures. The associated temporal kernel has a pure exponential shape, and its characteristic correlation time scale is $T_{\mathbf{k}}$. This is reminiscent of a dynamics for Fourier modes similar to the one of a Ornstein-Uhlenbeck process, as early recognized by Komorowski & Peszat (2004) and Eyink & Benveniste (2013). We intend to model the turbulent velocity field, up to second-order statistics, in the asymptotic regime of infinite Reynolds numbers. In a Gaussian framework, a given spatio-temporal covariance function of velocity components can be achieved by a stochastic process stirred by Gaussian space-time white noise. The very purpose of the present article is the formulation and simulation of such an incompressible random velocity field, where viscous corrections will be furthermore included in order to be compared to experimental measurements and DNSs of the Navier-Stokes equations. To do so, we not only need to impose viscous corrections to the spatial formulation, but also come up with a proper temporal dynamics that would ensure a consistent evolution of Fourier modes, giving rise to a smooth behavior at any wave vector, while respecting the causal nature of the underlying time marching. This will be achieved using the multi-layered evolution proposed by Viggiano et al. (2020). Also, each step of the formulation will be compared to DNSs data, in both space and time. The numerical data provided by the Johns Hopkins Turbulence Database (Li et al., 2008) allow such a comparison. Preliminary analysis of this DNS database has been performed by Reneuve (2019), and we will reproduce them here for convenience and completeness.
@@
As already mentioned, the purpose of the present article is the formulation and simulation of a Gaussian stochastic representation of the spatio-temporal structure of turbulence, which corresponds to the reproduction of the second-order statistical signatures that already encompass several key properties. Nonetheless, it is well known that the turbulent velocity field is non-Gaussian, as can be seen by the behavior through scales of higher-order statistical quantities, such as third-order moment of velocity increments that quantifies the asymmetry of Probability density Functions (PDFs) and energy transfers, and the fourth-order one which focuses on the heavy-tail nature of these PDFs (Frisch, 1995). These non-Gaussian features of the velocity field are referred to as the intermittency phenomenon which has a strong connection to multifractality. Several spatial random fields have been already proposed in the literature that are able to reproduce many aspects of this phenomenon, in particular the distributional nature of the dissipation field (Yaglom, 1962, 1966; Mandelbrot, 1972, 1974; Kahane & Peyrière, 1976; Kahane, 1985; Meneveau & Sreenivasan, 1987, 1991; Rhodes & Vargas, 2014), and their extension to the velocity field (Juneja et al., 1994; Biferale et al., 1998; Arneodo et al., 1998; Bacry et al., 2001; Robert & Vargas, 2008; Chevillard et al., 2010; Bacry et al., 2012; Pereira et al., 2016; Granero-Belinchón et al., 2018; Chevillard et al., 2019; Muzy, 2019; Chevillard et al., 2020), with generalizations towards structures and turbulent magnetic fields (Rosales & Meneveau, 2006, 2008; Durrive et al., 2020, 2022; Lübke et al., 2024). Let us mention that these aforementioned models are usually time-independent and they are meant to give a stochastic representation of statistically stationary solution of the Navier-Stokes equations. Recent works provide some generalization of these random fields with a given time evolution (Reneuve & Chevillard, 2020; Peixoto Considera & Thalabard, 2023). We keep the generalization of the present spatio-temporal random field for future works, which will be presented in a second article (in preparation).
@@
This article is organized as follows. In Chapter 2, we recall several important ingredients of the phenomenology of fluid turbulence based on second-order statistical quantities. Focusing on the spatial structure, we present in Section 2.1 basic statistical properties such as the viscous independence of velocity variance and dissipation per unit of mass, the wave-vector dependence of the power spectral density and the respective scale dependence of the spatial structure functions. This allows us to define with precision several key parameters of the formulation, such as the mean Hölder exponent $H=1/3$, the Kolmogorov universal constant $C_{2}$ and the dissipative length scale $\eta_{K}$. These predictions are reformulated for a Gaussian model and compared with a DNS in Section 2.2. We then discuss in Section 2.3 the behavior of characteristic time scales. In particular, we explore the implication of the sweeping effect, which has for statistical consequence that the characteristic turnover time scale of eddies of a given size is independent of the scale and coincides with the large integral time scale. Following these observations, we deduce the implied statistical structure of the velocity field based on the second-order temporal structure function and the associated temporal PSD in Section 2.4. We then present our spatio-temporal model in Chapter 3. In Section 3.1, we begin by reformulating former considerations with periodic boundary conditions, allowing to unambiguously consider Fourier modes since they are well defined in this context (unlike the former continuous formulation in full space). We formulate the model in full generality, for any Hölder exponent $0<H<1$ and for abitrary dependence of the characteristic time $T_{\mathbf{k}}$ on the wave-vector amplitude $|\mathbf{k}|$. In order to achieve the statistical temporal structure for the description of turbulence, we propose in Section 3.2 a causal stochastic evolution capable of generating such a field starting from any initial conditions. To do so, we first need to define the Fourier transform of a white noise, a formulation that we provide in subsection 3.2.1. This eventually allows us to define in subsection 3.2.2 a Gaussian random field that mimics the spatial structure depicted in Sections 2.1 and 3.1, which can be seen as a fractional Gaussian vector field. In subsection 3.2.3, we recall the associated stochastic dynamics of Chaves et al. (2003), and generalize it to a smooth and causal context, even toward an infinitely differentiable setup, in subsection 3.2.4. An effective numerical schemes is proposed in Chapter 4. We gather conclusions and perspectives in Chapter 5. Throughout the presentation of the model, we provide a precise comparison between the theoretical predictions of the model with both its numerical representations and the DNS data extracted from the Johns Hopkins turbulence database (Li et al., 2008). These comparisons concern both the spatial and temporal structure of velocity fields. We also provide key technical developments in the appendices.
@@
Let us consider the velocity vector field $\mathbf{u}^{\nu}(t,\mathbf{x})\in\mathbb{R}^{3}$ for $\mathbf{x}\in\mathbb{R}^{3}$ and $t\geq 0$, which reads component-wise $\mathbf{u}^{\nu}=(u_{i}^{\nu})_{1\leq i\leq 3}$, a solution of the Navier-Stokes equations, for incompressible fluids of unit density,
@@
$\sigma^{2}=\lim_{\nu\to 0}\lim_{t\to\infty}\mathbb{E}\left[|\mathbf{u}^{\nu}(t,\mathbf{x})|^{2}\right],$ (2)
@@
are positive numbers. Starting from the randomly forced Navier-Stokes equations (1), the expectations in (2) and (3) are taken over the the forcing realisations. From a practical point of view, we will assume that expectations can be accurately replaced by empirical averages over space and/or time, which is in particular guaranteed by the statistical homogeneity and stationarity observed in DNSs and experiments. In the asymptotic limit of infinite Reynolds numbers, i.e. in the vanishing viscosity limit $\nu\to 0$, the finite quantities $\sigma$ (2) and $\varepsilon$ (3) are expected to be related according to the dimensional prediction
@@
To characterize this behavior, usually referred to as the cascade phenomenon, the power spectral density (PSD) $E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})$ has been measured and estimated in many studies. This quantity is defined to be the Fourier transform of the trace of the covariance tensor of the velocity vector field, that is - using Einstein’s summation convention over repeated indices -
@@
Ìt can be checked that $E_{\nu}^{\text{\tiny{E,long}}}(k)$ is a positive and even function of $k$. The three-dimensional PSD $E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})$ (5), which is a function of the norm $k=|\mathbf{k}|$ of the wave vector $\mathbf{k}$, and the one-dimensional longitudinal PSD $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (8) are related as follows:
@@
Going back to the aforementioned phenomenological aspects of fluid turbulence, a manifestation of the underlying cascading process of energy can be read on the PSD, which will follow, in the limit of infinite Reynolds numbers, a power-law behavior at large wave numbers. This is known as the Kolmogorov $k^{-5/3}$” spectrum (Frisch, 1995). At a given finite Reynolds number, a realistic parameterization of the longitudinal PSD $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (8) would require a regularization at small wave numbers of order $1/L$, i.e. around the inverse integral length scale stemming from the forcing term, and a dissipative cut-off at large wave numbers $1/\eta_{K}$, where $\eta_{K}=O\left((\nu^{3}/\varepsilon)^{1/4}\right)$ is known as Kolmogorov’s dissipative length scale. Sophisticated models that do so have been proposed in the literature (Pope, 2000; Meyers & Meneveau, 2008). For the sake of simplicity, we will be working with the following simple but realistic parameterization:
@@
which is of order of the Kolmogorov dissipative length scale $\eta_{K}$. $\Gamma\left(z\right)$ denotes here the gamma function. The expression of the pre-factor in the definition of $\eta_{d}$ (12) is derived below. The so-called Kolmogorov constant $C_{2}$ has been estimated on experimental data and is close to $C_{2}\approx 2$, see Pope (2000). From the proposed expressionl for $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (10), we deduce from the identity (9) the expression of the three-dimensional PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ (5), which has a particularly simple expression in the vanishing viscosity limit, i.e.
@@
$\sigma^{2}=\lim_{\nu\to 0}\sigma_{\nu}^{2}=6D_{2}\int_{k=0}^{\infty}|\mathbf{k}|_{1/L}^{-5/3}\ {\mathrm{d}}\mathbf{k},$ (15)
@@
$\delta_{\ell}^{\text{\tiny{long}}}\mathbf{u}^{\nu}(t,\mathbf{x})=\left(\mathbf{u}^{\nu}(t,\mathbf{x}+\ell)-\mathbf{u}^{\nu}(t,\mathbf{x})\right)\cdot\frac{\ell}{|\ell|}\ ,$ (19)
@@
$=2\int_{k\in\mathbb{R}}\left(1-e^{2i\pi k|\ell|}\right)E^{\text{\tiny{E,long}}}(k){\mathrm{d}}k=4D_{2}\int_{k=0}^{\infty}\left(1-\cos(2\pi k|\ell|)\right)|k|_{1/L}^{-5/3}{\mathrm{d}}k$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}4D_{2}|\ell|^{2/3}\int_{k=0}^{\infty}\left(1-\cos(2\pi k)\right)k^{-5/3}{\mathrm{d}}k=3D_{2}(2\pi)^{2/3}\Gamma(1/3)|\ell|^{2/3}.$ (21)
As can be seen from the behavior of $S_{2}^{\text{\tiny{long}}}(\ell)$ (2.1) on small scales $|\mathbf{\ell}|\to 0$, the second-order structure function exhibits a power-law behavior with respect to the scale $\ell$ with the exponent $2/3$. Using then the expression of $D_{2}$ (11), and expressing the remaining multiplicative constant in units of the average dissipation $\varepsilon$ (18), we finally get
$S_{2}^{\text{\tiny{long}}}(\ell)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}C_{2}|\varepsilon\ell|^{2/3},$ (22)
@@
$\delta_{\ell}u_{i}^{\nu}(t,\mathbf{x})=u_{i}^{\nu}(t,\mathbf{x}+\ell)-u_{i}^{\nu}(t,\mathbf{x}),$ (23)
@@
$S_{2,ij}(\ell)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}C_{2}|\varepsilon\ell|^{2/3}\left[-\frac{1}{3}\frac{\ell_{i}\ell_{j}}{|\ell|^{2}}+\frac{4}{3}\delta_{ij}\right].$ (25)
@@
$S_{2}(\ell)\equiv S_{2,ii}(\ell)=\lim_{\nu\to 0}\mathbb{E}\left[\left|\delta_{\ell}\mathbf{u}^{\nu}\right|^{2}\right]$ $=2\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\ell}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}\frac{11}{3}C_{2}|\varepsilon\ell|^{2/3}.$ (26)

### v2

# A spatio-temporal random synthetic turbulent velocity field: The underlying Gaussian structure

@@ Abstract
We develop, simulate and extend an initial proposition by Chaves et al. (2003) concerning a random incompressible vector field able to reproduce key ingredients of three-dimensional turbulence in both space and time. In this article, we focus on the important underlying Gaussian framework. Presently, the statistical spatial structure of this velocity field is consistent with a divergence-free fractional Gaussian vector field that encodes all known properties of homogeneous and isotropic fluid turbulence at a given finite Reynolds number, up to second-order statistics. The temporal structure of the velocity field is introduced through a stochastic evolution of the respective Fourier modes. In the simplest picture, Fourier modes evolve according to an Ornstein-Uhlenbeck process, where the characteristic time scale depends on the wave-vector amplitude. For consistency with direct numerical simulations (DNSs) of the Navier-Stokes equations, this time scale is inversely proportional to the wave vector amplitude. As a consequence, the characteristic velocity that governs the eddies is independent of their size and is related to the velocity standard deviation, which is consistent with some features of the so-called sweeping effect. To ensure differentiability in time while respecting the Markovian nature of the evolution, we use the methodology developed by Viggiano et al. (2020) to propose a fully consistent stochastic picture. We finally derive analytically all statistical quantities in a continuous setup and develop precise and efficient numerical schemes of the corresponding periodic framework. Both exact predictions and numerical estimations of the model are compared to DNSs provided by the Johns Hopkins database.
@@ 1 Introduction
Fully developed fluid turbulence is emblematic of fluid mechanics and out-of-equilibrium statistical physics. It is usually observed in various natural flows and studied, in simpler situations than those of geophysical flows, in laboratory experiments and numerical simulations of the forced Navier-Stokes equations. An ancient tradition of measurements and observations has allowed the development of a very precise and self-consistent phenomenology capable of describing its statistical structure with few free parameters (Monin & Yaglom, 1971; Tennekes & Lumley, 1972; Frisch, 1995; Pope, 2000). From the theoretical side, the velocity of a turbulent fluid can be seen as the vanishing viscosity limit of the statistically stationary, isotropic, and homogeneous solution of the incompressible Navier-Stokes equations sustained by a random (Eswaran & Pope, 1988), possibly deterministic (Vincent & Meneguzzi, 1991), force. Because of the interplay between non-linear and non-local aspects of this dynamics, with non-trivial influence of viscosity, the rigorous study of these solutions remains a fantastic mathematical challenge. The main purpose of this article is the presentation, extension, and simulation of a spatio-temporal turbulent velocity which Fourier modes evolve according to a Markovian dynamics in the simplest situation, supplemented by additional underlying layers in a way that we will make precise in a more sophisticated version, able to reproduce several key statistical signatures. This stochastic model encompasses the spatial multi-scale nature of the observed turbulent velocity field, and a associated random evolution in time. In the infinite Reynolds number limit and when the analysis is restricted to second-order statistical laws, as captured in particular by a Gaussian framework, and assuming the model to evolve according to a single-layer Markovian dynamics, the present approach becomes equivalent to that proposed by Chaves et al. (2003). As far as we know, such models are unrelated to Navier-Stokes solutions, and we will therefore consider them as being synthetic, although they reflect important aspects of fluid turbulence.
@@
The purpose of the present article is to propose a random incompressible field that evolves according to a stochastic evolution consistent with the temporal structure of turbulence, which has been repeatedly observed, mostly in DNSs (Kaneda et al., 1999; Chevillard et al., 2005; Favier et al., 2010; Canet et al., 2017; Gorbunova et al., 2021). It is more difficult to observe this evolution in experimental measurements because it requires the absence of a mean flow, which is present in wind tunnels and jets. However, Gorce & Falcon (2022) have conducted measurements of the temporal evolution of a flow stirred by magnetic particles able to generate turbulence without a mean flow, and temporal statistics were found to be similar to those observed in aforementioned numerical simulations. The temporal spectrum, defined as the variance of the temporal Fourier modes at a given frequency $\omega$, behaves in a universal fashion as a power-law $\omega^{-5/3}$”, with thus a similar exponent as observed on the spatial spectrum. Although understood and predicted by Tennekes (1975), this power-law decay can be seen as counterintuitive and differs drastically from similarity arguments which would predict a much faster decrease, i.e. with a 2-2 power-law exponent, as has been observed in the Lagrangian framework, i.e. following fluid particles along their trajectory (Yeung, 2002; Toschi & Bodenschatz, 2009; Pinton & Sawford, 2012). Temporal regularity of the velocity field is imposed by this power-law, which is thus the same as the spatial one, and turns out to have important consequences. In fact, the characteristic velocity of eddies of a given size is independent of that scale and turns out to be the one at large scale: this is the so-called sweeping effect.
From a statistical viewpoint, both this power-law decay of the temporal power spectral density (PSD) and the scale-independent characteristic velocity of eddies can be included in the model by setting a decorrelation time scale $T_{\mathbf{k}}$ of Fourier modes proportional to the inverse wave-vector amplitude and whose multiplicative constant is inversely proportional to the velocity standard deviation allowing a consistent physical picture and proper dimension. This time scale dependent on $k$ is at the core of the random picture of Chaves et al. (2003) (see also earlier propositions of Holzer & Siggia (1994), Komorowski & Papanicolaou (1997) and Fannjiang & Komorowski (2000), and related discussions by Komorowski & Peszat (2004) and Eyink & Benveniste (2013)). Let us mention that, from a stochastic modeling point of view, building up a random field able to reproduce the observed aforementioned temporal structure is far from being obvious. In particular, it is known from earlier investigations by Mandelbrot & Van Ness (1968) that fractional regularity usually requires a non Markovian dynamics and calls for sophisticated formulations of the fractional Brownian motions, as proposed by Alòs et al. (2000) and latterly by Chevillard (2017). To this regard, the proposition of Chaves et al. (2003) is very original, and shows that one could build instead the peculiar temporal regularity of the turbulent velocity in a Markovian way using the randomness introduced in the spatial dimensions using respective Fourier modes.
@@
Nonetheless, it is well known that the turbulent velocity field is non-Gaussian, as can be seen by the behavior through scales of higher-order statistical quantities, such as third-order moment of velocity increments that quantifies the asymmetry of Probability density Functions (PDFs) and energy transfers, and the fourth-order one which focuses on the heavy-tail nature of these PDFs (Frisch, 1995). These non-Gaussian features of the velocity field are referred to as the intermittency phenomenon which has a strong connection to multifractality. Several spatial random fields have been already proposed in the literature that are able to reproduce many aspects of this phenomenon, in particular the distributional nature of the dissipation field (Yaglom, 1962, 1966; Mandelbrot, 1972, 1974; Kahane & Peyrière, 1976; Kahane, 1985; Meneveau & Sreenivasan, 1987, 1991; Rhodes & Vargas, 2014), and their extension to the velocity field (Juneja et al., 1994; Biferale et al., 1998; Arneodo et al., 1998; Bacry et al., 2001; Robert & Vargas, 2008; Chevillard et al., 2010; Bacry et al., 2012; Pereira et al., 2016; Granero-Belinchón et al., 2018; Chevillard et al., 2019; Muzy, 2019; Chevillard et al., 2020), with generalizations towards structures and turbulent magnetic fields (Rosales & Meneveau, 2006, 2008; Durrive et al., 2020, 2022; Lübke et al., 2024) and solid-state physics (Lakhal et al., 2025). Let us mention that these aforementioned models are usually time-independent and they are meant to give a stochastic representation of statistically stationary solution of the Navier-Stokes equations. Recent works provide some generalization of these random fields with a given time evolution (Reneuve & Chevillard, 2020; Peixoto Considera & Thalabard, 2023). We keep the generalization of the present spatio-temporal random field to a multifractal framework for future works.
@@
This article is organized as follows. In Chapter 2, we recall several important ingredients of the phenomenology of fluid turbulence based on second-order statistical quantities. Focusing on the spatial structure, we present in Section 2.1 basic statistical properties such as the viscosity independence of velocity variance and dissipation per unit of mass, the wave-vector dependence of the power spectral density and the respective scale dependence of the spatial structure functions. This allows us to define with precision several key parameters of the formulation, such as the mean Hölder exponent $H=1/3$, the Kolmogorov universal constant $C_{2}$ and the dissipative length scale $\eta_{K}$. These predictions are reformulated for a Gaussian model and compared with a DNS in Section 2.2. We then discuss in Section 2.3 the behavior of characteristic time scales. In particular, we explore the implication of the sweeping effect, which has for statistical consequence that the characteristic turnover time scale of eddies of a given size is independent of the scale and coincides with the large integral time scale. Following these observations, we deduce the implied statistical structure of the velocity field based on the second-order temporal structure function and the associated temporal PSD in Section 2.4. We then present our spatio-temporal model in Chapter 3. In Section 3.1, we begin by reformulating former considerations with periodic boundary conditions. In order to achieve the statistical temporal structure for the description of turbulence, we propose in Section 3.2 a Markovian stochastic evolution capable of generating such a field starting from any initial conditions. In subsection 3.2.1 we define a Gaussian random field that mimics the spatial structure depicted in Sections 2.1 and 3.1, which can be seen as a fractional Gaussian vector field. In subsection 3.2.2, we recall the associated stochastic dynamics of Chaves et al. (2003), and generalize it to a smooth context, even toward an infinitely differentiable setup, in subsection 3.2.3. This will be achieved using the multi-layered evolution proposed by Viggiano et al. (2020). An effective numerical schemes is proposed in Chapter 4. We gather conclusions and perspectives in Chapter 5. Also, each step of the formulation will be compared to DNSs data, in both space and time. The numerical data provided by the Johns Hopkins Turbulence Database (Li et al., 2008) allow such a comparison. Preliminary analysis of this DNS database has been performed by Reneuve (2019), and we will reproduce them here for convenience and completeness. We also provide key technical developments in the appendices.
@@
Let us consider the velocity vector field $\mathbf{u}^{\nu}(t,\mathbf{x})\in\mathbb{R}^{3}$ for $\mathbf{x}\in\mathbb{R}^{3}$ and $t\geq 0$, which reads component-wise $\mathbf{u}^{\nu}=(u_{i}^{\nu})_{1\leq i\leq 3}$, solution of the Navier-Stokes equations, for incompressible fluids of unit density,
@@
$\sigma^{2}=\lim_{\nu\to 0}\mathbb{E}\left[|\mathbf{u}^{\nu}|^{2}\right],$ (2)
@@
are positive numbers. Starting from the randomly forced Navier-Stokes equations (1), the expectations in (2) and (3) are taken over the the forcing realisations. From a practical point of view, we will assume that expectations can be accurately replaced by empirical averages over space and/or time, which is in particular guaranteed by the statistical homogeneity and stationarity observed in DNSs and experiments. In the asymptotic limit of infinite Reynolds numbers, i.e. in the vanishing viscosity limit $\nu\to 0$, the finite quantities $\sigma$ (2) and $\varepsilon$ (3) are expected to be related according to the dimensional prediction (see the classical textbooks by Frisch (1995) and Pope (2000) for instance)
@@
To characterize this behavior, usually referred to as the cascade phenomenon, the PSD $E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})$ has been measured and estimated in many studies. This quantity is defined to be the Fourier transform of the trace of the covariance tensor of the velocity vector field, that is - using Einstein’s summation convention over repeated indices -
@@
It can be checked that $E_{\nu}^{\text{\tiny{E,long}}}(k)$ is a positive and even function of $k$. The three-dimensional PSD $E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})$ (5), which is a function of the norm $k=|\mathbf{k}|$ of the wave vector $\mathbf{k}$, and the one-dimensional longitudinal PSD $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (8) are related as follows:
@@
Going back to the aforementioned phenomenological aspects of fluid turbulence, a manifestation of the underlying cascading process of energy can be read on the PSD, which will follow, in the limit of infinite Reynolds numbers, a power-law behavior at large wave numbers. This is known as the Kolmogorov $k^{-5/3}$” spectrum (Frisch, 1995). At a given finite Reynolds number, a realistic parameterization of the longitudinal PSD $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (8) would require a regularization at small wave numbers of order $1/L$, i.e. around the inverse integral length scale stemming from the forcing term, and a dissipative cut-off at large wave numbers $1/\eta_{K}$, where $\eta_{K}\propto(\nu^{3}/\varepsilon)^{1/4}$is known as Kolmogorov’s dissipative length scale. Sophisticated models that do so have been proposed in the literature (Pope, 2000; Meyers & Meneveau, 2008). For the sake of simplicity, we will be working with the following simple but realistic parameterization:
@@
which is of order of the Kolmogorov dissipative length scale $\eta_{K}$. $\Gamma\left(z\right)$ denotes here the gamma function. The expression of the pre-factor in the definition of $\eta_{d}$ (12) is derived below, and is chosen such that to obtain ultimately (22). The so-called Kolmogorov constant $C_{2}$ has been estimated on experimental data and is close to $C_{2}\approx 2$, see Pope (2000). From the proposed expression for $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (10), we deduce from the identity (9) the expression of the three-dimensional PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ (5), which has a particularly simple expression in the vanishing viscosity limit, i.e.
@@
$\sigma^{2}=\lim_{\nu\to 0}\sigma_{\nu}^{2}=6D_{2}\int_{k=0}^{\infty}|k|_{1/L}^{-5/3}\ {\mathrm{d}}k,$ (15)
@@
$\delta_{\ell}^{\text{\tiny{long}}}\mathbf{u}^{\nu}(t,\mathbf{x})=\left(\mathbf{u}^{\nu}(t,\mathbf{x}+\boldsymbol{\ell})-\mathbf{u}^{\nu}(t,\mathbf{x})\right)\cdot\frac{\boldsymbol{\ell}}{\ell}\ ,$ (19)
@@
$=2\int_{k\in\mathbb{R}}\left(1-e^{2i\pi k\ell}\right)E^{\text{\tiny{E,long}}}(k){\mathrm{d}}k=4D_{2}\int_{k=0}^{\infty}\left(1-\cos(2\pi k\ell)\right)|k|_{1/L}^{-5/3}{\mathrm{d}}k$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}4D_{2}\ell^{2/3}\int_{k=0}^{\infty}\left(1-\cos(2\pi k)\right)k^{-5/3}{\mathrm{d}}k=3D_{2}(2\pi)^{2/3}\Gamma(1/3)\ell^{2/3}.$ (21)
As can be seen from the behavior of $S_{2}^{\text{\tiny{long}}}(\ell)$ (2.1) on small scales $\ell\to 0$, the second-order structure function exhibits a power-law behavior with respect to the scale $\ell$ with the exponent $2/3$. Using then the expression of $D_{2}$ (11), and expressing the remaining multiplicative constant in units of the average dissipation $\varepsilon$ (18), we finally get
$S_{2}^{\text{\tiny{long}}}(\ell)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}C_{2}|\varepsilon\ell|^{2/3},$ (22)
@@
$\delta_{\ell}u_{i}^{\nu}(t,\mathbf{x})=u_{i}^{\nu}(t,\mathbf{x}+\boldsymbol{\ell})-u_{i}^{\nu}(t,\mathbf{x}),$ (23)
@@
$S_{2,ij}(\ell)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}C_{2}|\varepsilon\ell|^{2/3}\left[-\frac{1}{3}\frac{\ell_{i}\ell_{j}}{\ell^{2}}+\frac{4}{3}\delta_{ij}\right].$ (25)
@@
$S_{2}(\ell)\equiv S_{2,ii}(\ell)=\lim_{\nu\to 0}\mathbb{E}\left[\left|\delta_{\ell}\mathbf{u}^{\nu}\right|^{2}\right]$ $=2\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}\frac{11}{3}C_{2}|\varepsilon\ell|^{2/3}.$ (26)